A fishing boat owner brings 500 fish to market and the market price is $4 per fish. Her average variable cost of 500 fish is $1 and the fixed cost of the boat is $2,000. What is her profit per fish?

Group of answer choices

A. $1.

B. -$1.

C. $5,000.

D. $25,000.

E. $500,000.
Answer #1

Answer

average fixed cost =total fixed cost /quantity

=2000/500

=4

average total cost =average varible cost +average fixed costs

=1+4=5

Profit per unit =price -average total costs

=4-5

-$1

the profit per fish is -$1

Option B